# TVBench: Redesigning Video-Language Evaluation

### v1

@@ Abstract
Large language models have demonstrated impressive performance when integrated with vision models even enabling video understanding. However, evaluating these video models presents its own unique challenges, for which several benchmarks have been proposed. In this paper, we show that the currently most used video-language benchmarks can be solved without requiring much temporal reasoning. We identified three main issues in existing datasets: (i) static information from single frames is often sufficient to solve the tasks (ii) the text of the questions and candidate answers is overly informative, allowing models to answer correctly without relying on any visual input (iii) world knowledge alone can answer many of the questions, making the benchmarks a test of knowledge replication rather than visual reasoning. In addition, we found that open-ended question-answering benchmarks for video understanding suffer from similar issues while the automatic evaluation process with LLMs is unreliable, making it an unsuitable alternative. As a solution, we propose TVBench, a novel open-source video multiple-choice question-answering benchmark, and demonstrate through extensive evaluations that it requires a high level of temporal understanding. Surprisingly, we find that most recent state-of-the-art video-language models perform similarly to random performance on TVBench, with only Gemini-Pro and Tarsier clearly surpassing this baseline.
@@ 1 Introduction
Vision Language Models have gained popularity, benefiting from both the progress made in natural language processing and the surge of foundation models for vision tasks with strong generalization capabilities. Recently, video-language models have been introduced (Xu et al., 2021; Lin et al., 2023; Zhang et al., 2023a; Xue et al., 2023), aiming to replicate the success achieved in the image domain. To evaluate their performance, visual question answering has emerged as a key task requiring both textual and visual reasoning. With the rapid model development and release cycles, having a reliable and robust benchmark is crucial in measuring progress and guiding research efforts.
There are two main approaches to designing question-answering benchmarks for videos: multiple-choice question answering (MCQA)  (Wang et al., 2024; Li et al., 2024) and open-ended question answering (Yu et al., 2019; Xu et al., 2017). The most popular and commonly used benchmark for MCQA is MVBench (Li et al., 2024), which is the most downloaded video dataset with over 104K monthly downloads as of September 2024 and the second most downloaded across both image and video visual question-answering benchmarks on HuggingFace1, despite being released only recently. As such, its reliability as a valid benchmark is of utmost importance. But how much video understanding does MVBench truly measure?
@@
    $\rightarrow$ To address this, we generate questions using text templates, preventing grammatical issues that could be exploited by text-only models.
@@
    $\rightarrow$ We design questions that can be answered purely from the video content, without relying on prior world knowledge.
As a result, TVBench measures the temporal understanding of video-language models in contrast to MVBench, as shown in Fig. 1. In this setting, text-only and single-frame models, such as Llama3 and GPT-4o, perform at random chance level on TVBench, despite achieving competitive results on MVBench. Surprisingly, even the most recent state-of-the-art video-language models perform close to random chance on TVBench, with Tarsier and Gemini 1.5 Pro clearly outperforming the random baseline. For these models, shuffling and reversing the videos lead to significant performance drops, unlike in MVBench, further highlighting the value of TVBench as a temporal video benchmark.
@@ 2 Related work
Multiple-choice question answering. CLEVRER (Yi et al., 2019) assesses reasoning about object interaction in synthetic videos. Perception Test (Patraucean et al., 2024) was introduced to evaluate visual perception in multimodal settings, mainly in indoor scenes. (Bagad et al., 2023) uses synthetic generated data to evaluate the temporal understanding of early video-language models. EgoSchema (Mangalam et al., 2024) focuses on MCQA involving long egocentric videos. Recently, VideoHallucer (Wang et al., 2024) was introduced as a first attempt to define a video-language benchmark specifically designed for hallucination detection. MVBench (Li et al., 2024) aims to evaluate temporal understanding by defining 20 dynamic tasks specifically designed to require temporal reasoning throughout the entire video. However, our experiments demonstrate that many of these tasks are highly spatial and textual biased, failing to evaluate temporal understanding effectively. We propose a new benchmark that requires a high level of spatiotemporal understanding across different tasks in order to be solved.
@@ 3 Problems in Video MCQA Benchmarks
In this section, we identify two key shortcomings in current video multiple-choice question-answering benchmarks, as demonstrated on MVBench. First, we show that this benchmark contains strong spatial bias, meaning that questions can be answered without temporal understanding. Secondly, we demonstrate that MVBench also contains a strong textual bias, as many questions can be answered without even looking at the visual input.
@@ 3.1 Does Time Matter?
Figure 2: Spatial bias of MVBench video-language benchmark. We show different tasks of the MVBench benchmark and observe that the question can be answered without requiring any temporal understanding. For quantitative results see Tab. 1.
@@
We analyze this spatial bias in MVBench using different state-of-the-art image and video-language models. In Tab. 1 we focus on four different tasks i) scene transition ii) fine-grained pose iii) episodic reasoning and iv) fine-grained action. To assess whether solving these tasks requires temporal understanding, we compare the performance of image-language models, which receive only a single image, with video-language models that process both original and shuffled videos.
The image-language models demonstrate strong performance across all four tasks in Tab. 1, surpassing random baselines and often matching the capabilities of state-of-the-art video-language models. This is unexpected, as task names like Fine-grained Action suggest a need for temporal understanding. For this fine-grained task, the image-model GPT-4o achieves 49%, which is even slightly better than the state-of-the-art model Tarsier, which scores 48.5%. Similarly, for the other three tasks. Overall, GPT-4o achieves an average accuracy across all 20 tasks of 47.8%, which is 20.5% higher than the random performance of 27.3% on MVBench. This indicates that a large portion of the benchmark is affected by a spatial bias.
Additionally, shuffling the videos has minimal impact on the performance of video-language models, indicating that temporal information is not necessary to solve these tasks. Note, as confirmed in Sec. 5.3, the Tarsier model shows a significant drop in performance when videos are shuffled for tasks that require temporal understanding. This problem goes beyond these four tasks as shown in Tab. 4, VideoChat2 and Tarsier achieve an average accuracy across all 20 MVBench tasks of 51.1% and 67.7%, respectively. Shuffling video frames causes a performance drop of 0.2% and 1.2%, respectively, indicating that the spatial bias affects not only the tasks analyzed in Tab. 1 but the entire dataset. Additionally, VideoChat2 provides the same answer for both the original and shuffled videos 91% of the time, while Tarsier shows agreement on 82% of the questions.
In Fig. 2 we show examples corresponding to tasks analyzed in Tab. 1. The first question suggests that temporal understanding is needed as it requires identifying moving objects at the end of the video. Yet, the question can be answered without temporal understanding as no red objects moving or stationary exist in the whole video. The second and third examples are from the Fine-grained Action and Fine-grained Pose tasks respectively. In both cases, the information provided from any frame is enough to correctly answer the question.
@@
Table 1: Spatial bias of the MVBench video-language benchmark. Our evaluation reveals that temporal understanding is not required for solving temporal tasks of MVBench such as fine-grained action. We find that tasks can be solved with a random frame, and shuffling videos does not impact the performance of current video-language models.
@@ 3.2 Does Vision Matter?
Video benchmarks must be designed to prevent questions from being answered solely through common sense reasoning. Modern LLMs possess strong reasoning skills, which can exploit the information within the question and candidate sets in MCQA video language evaluation benchmarks. This introduces substantial textual bias in the results, as models might correctly answer questions without leveraging the video content.
To examine the textual bias, we evaluate the performance of text-only LLMs on MVBench, as shown in Tab. 2. Our findings show that LLMs can eliminate incompatible candidates easily, greatly outperforming a random model. A text-only model achieves even competitive results across four tasks with video-language models without using video input. For example, Llama 3 performs nearly on par with 44.5% compared to the video-language model Tarsier, which achieves 46.5% on the Action Count task. This goes beyond the four tasks, as Llama 3 achieves an average performance across all 20 tasks of 38.1%, which is 10.8% higher than the random chance baseline of 27.3%. We have identified three key sources of this textual bias.
Bias from LLM-based QA generation. Collecting and manually annotating large datasets for training and evaluation is very costly. Automatic and semi-automatic collection and annotation processes are commonly used to address this challenge (Li et al., 2024; Mangalam et al., 2024). This includes techniques such as automatic QA pair generation with LLMs. ChatGPT plays a fundamental role in QA generation for 11 of the 20 tasks in the MVBench dataset. However, this introduces unrealistic candidates and QA pairs with excessive information. Fig. 3 presents examples of QA pairs that can be resolved merely with text information. Questions 1 and 2 belong to the Action Antonym task, where an LLM is prompted to generate the antonym of the actual action shown in the video. The answers generated are either unrealistic, as one cannot “remove something into something,” or consistently incorrect, such as “not sure.”
@@
Bias from unbalanced sets. Unbalanced QA sets also hinder a robust evaluation process. For instance, the correct answer for the Action Count task on MVBench is ‘3’ for 90 out of 200 questions, while ‘9’ is only the correct answer for one question. A model with a similar bias might get higher results than random by chance. We have observed in our experiments that some text-only models such as GPT-4o have the same bias, predicting ‘3’ for 88 out of 200 samples. This phenomenon makes GPT-4o perform on par with the best video model with an accuracy of 44.0% and 46.5% respectively.
@@
Table 2: Textual bias of the MVBench video-language benchmark. Our analysis reveals that vision is not required for solving tasks from MVBench as text-only LLMs score high above the random baseline and nearly on par with video models.
@@ 4 Open-ended QA to the rescue?
Moreover, as shown in Tab. 3 open-ended QA does not solve the main issues of MCQA. The performance of text-only models is surprisingly strong; state-of-the-art LLMs can guess the answer solely from the question text for a significant number of questions, even without a candidate list. This includes questions such as Which hand of the person in black wears a watch? or What color is the pants of a person wearing black clothes?, which correct answers are Left hand and Black. The first question can be answered just with prior knowledge as people commonly wear the watch on the left hand, while in the second one, the question gives too much information, the person is wearing black clothes. Similar to the findings for MCQA on spatial bias in Sec. 3.1, when using a single random frame for image-text models such as GPT-4o, performance reaches 60.6% and 46.4%, approaching the video-language model’s 80.3% and 61.6%, respectively. In addition, the performance of Tarsier-34B does not significantly drop—on average less than 3%—when the input videos are shuffled, indicating the low temporal understanding required for solving the benchmarks. These results show that open-ended video-language benchmarks also exhibit strong spatial bias, not requiring temporal understanding to be solved.
@@ 5 TVBench: A temporal video question answering Benchmark
MSVD-QA Llama3 70B text-only 29.1 2.6 51.9 2.8 +22.8
GPT-4o text-only 29.0 2.5 44.2 2.4 +15.2
GPT-4o single image 60.6 3.6 75.3 3.8 +14.7
@@
MSRVTT-QA Lama3 70B text-only 23.9 2.4 47.8 2.6 +23.9
GPT-4o text-only 23.3 2.3 42.4 2.3 +19.1
GPT-4o single image 34.2 2.7 50.8 2.7 +16.6
@@
ActivityNet-QA Lama3 70B text-only 25.3 2.6 32.6 1.9 +7.3
GPT-4o text-only 27.1 2.5 33.9 1.8 +6.8
GPT-4o single image 46.4 3.2 56.2 2.9 +9.8
@@
We propose TVBench, a new benchmark specifically created to evaluate temporal understanding in video QA. We adopt a multiple-choice QA approach to prevent the problems described in Sec. 4 regarding open-ended VQA evaluation. The main design principles of TVBench are derived from, and address, the problems listed in Sec. 3. These principles can be applied to any video QA benchmark to make it a temporally demanding benchmark. Tab. 5 in Appendix A provides an overview of the tasks, questions, and answers candidates used in our benchmark.
5.1 Developing a temporal video-language benchmark
This section explains the key solutions implemented in TVBench to address the issues identified in Sec. 3 of current video MCQA evaluation benchmarks.
Defining hard answer candidates. To address Problem 1, the temporal constraints in the question should be essential for arriving at the correct answer. Ignoring these constraints should result in incorrect or random responses. This involves designing time-sensitive questions and challenging candidates that cannot be easily disregarded without temporal understanding, which is essential for accurately evaluating the capabilities of current video-language models. Fig. 5 illustrates our main approach to solving Problem 1. Instead of including random, easy negative candidates, we define hard candidates that cannot be discarded without temporal information. Consequently, single image models and video models fed with shuffled variations of the original video should achieve random performance. We defined 10 temporally challenging tasks that either require repetition counting (Action Count), properties about moving objects (Object Shuffle, Object Count, Moving Direction), temporal localization (Action Localization, Unexpected Action), temporal sequential ordering (Action Sequence, Scene Transition, Egocentric Sequence) and distinguishing between temporally hard Action Antonyms such as Standing up and Sitting down.
Define questions that are not overly informative. Contrary to LLM-based generation, we apply straightforward templates to mitigate the effect of text-biased QA pairs, mitigating Problem 2. Tab 5 shows how questions and candidates do not provide enough information to give the correct answer solely from text. For most tasks, we keep the same candidate set within the same task. This includes Action Count, Object Count, Object Shuffle, Action Localization, Unexpected Action, and Moving Direction, with each candidate being the correct answer for an equal number of questions, ensuring a balanced dataset. This approach allows us to reduce textual bias while keeping the complexity of the visual task. For instance, the model still needs to recognize the unexpected action to localize it, even though the text does not refer to that action. There are always two candidates for Action Sequence and Scene Transition: two actions from the video and two possible sequences of scenes, respectively. Finally, the candidate set for Egocentric Sequence always contains the correct order of actions and random variations changing the order as negative candidates.
@@ 5.1 Developing a temporal video-language benchmark
Omitting questions that require prior knowledge. Solving Problem 2, mitigating the overreliance on world knowledge, requires providing questions and candidates that contain only the necessary information, specifically removing factual information that the LLM can exploit. Pre-existing knowledge should not influence the reasoning process, as this could lead to evaluating a different task than visual understanding. It becomes essential for the model to focus on the provided visual information rather than relying on common knowledge to find the correct answer. We remove from our proposal tasks such as Episodic Reasoning from MVBench, that are based on QA pairs about TV shows or movies. These samples cannot be corrected without manual intervention.

### v2

@@ Abstract
Large language models have demonstrated impressive performance when integrated with vision models even enabling video understanding. However, evaluating these video models presents its own unique challenges, for which several benchmarks have been proposed. In this paper, we show that the currently most used video-language benchmarks can be solved without requiring much temporal reasoning. We identified three main issues in existing datasets: (i) static information from single frames is often sufficient to solve the tasks (ii) the text of the questions and candidate answers is overly informative, allowing models to answer correctly without relying on any visual input (iii) world knowledge alone can answer many of the questions, making the benchmarks a test of knowledge replication rather than visual reasoning. In addition, we found that open-ended question-answering benchmarks for video understanding suffer from similar issues while the automatic evaluation process with LLMs is unreliable, making it an unsuitable alternative. As a solution, we propose TVBench, a novel open-source video multiple-choice question-answering benchmark, and demonstrate through extensive evaluations that it requires a high level of temporal understanding. Surprisingly, we find that most recent state-of-the-art video-language models perform similarly to random performance on TVBench, with only a few models such as Qwen2-VL, and Tarsier clearly surpassing this baseline.
@@ 1 Introduction
Vision Language Models have gained popularity, benefiting from both the progress made in natural language processing and the surge of foundation models for vision tasks with strong generalization capabilities. Recently, video-language models have been introduced (Xu et al., 2021; Lin et al., 2023; Xue et al., 2023), aiming to replicate the success achieved in the image domain. To evaluate their performance, visual question answering has emerged as a key task requiring both textual and visual reasoning. With the rapid model development and release cycles, having a reliable and robust benchmark is crucial in measuring progress and guiding research efforts.
There are two main approaches to designing question-answering benchmarks for videos: multiple-choice question answering (MCQA)  (Wang et al., 2024c; Li et al., 2024b) and open-ended question answering (Yu et al., 2019; Xu et al., 2017). The most popular and commonly used benchmark for MCQA is MVBench (Li et al., 2024b), which is the most downloaded video dataset with over 104K monthly downloads as of September 2024 and the second most downloaded across both image and video visual question-answering benchmarks on Hugging Face despite being released only recently. As such, its reliability as a valid benchmark is of utmost importance. But how much video understanding does MVBench truly measure?
@@
    $\rightarrow$ We design task-specific templates to generate questions that are not overly informative such that they cannot be answered solely by text.
@@
    $\rightarrow$ We design questions that can only be answered from the video content, without relying on prior world knowledge.
As a result, TVBench measures the temporal understanding of video-language models in contrast to MVBench, as shown in Fig. 1. In this setting, text-only and single-frame models, such as Gemini 1.5 Pro and GPT-4o, perform at random chance levels on TVBench, despite achieving competitive results on MVBench. Surprisingly, even recent state-of-the-art video-language models perform close to random chance on TVBench, with only a few models such as Qwen2-VL and Tarsier outperforming the random baseline. For these models, shuffling and reversing the videos lead to significant performance drops, unlike in MVBench, further verifying TVBench as a temporal video benchmark.
@@ 2 Related work
Multiple-choice question answering. CLEVRER (Yi et al., 2019) assesses reasoning about object interaction in synthetic videos. Perception Test (Patraucean et al., 2024) was introduced to evaluate visual perception in multimodal settings, mainly in indoor scenes. (Bagad et al., 2023) uses synthetic generated data to evaluate the temporal understanding of early video-language models. EgoSchema (Mangalam et al., 2024) focuses on MCQA involving long egocentric videos. Recently, VideoHallucer (Wang et al., 2024c) was introduced as a first attempt to define a video-language benchmark specifically designed for hallucination detection. MVBench (Li et al., 2024b) aims to evaluate temporal understanding by defining 20 dynamic tasks specifically designed to require temporal reasoning throughout the entire video. However, our experiments demonstrate that many of these tasks are highly spatial and textual biased, failing to evaluate temporal understanding effectively. We propose a new benchmark that requires a high level of spatiotemporal understanding across different tasks in order to be solved.
@@ 3 Problems in Video MCQA Benchmarks
In this section, we identify two key shortcomings in current video multiple-choice question-answering benchmarks, as demonstrated on MVBench. First, we show that this benchmark contains strong spatial bias, meaning that questions can be answered without temporal understanding. Secondly, we demonstrate that MVBench also contains a strong textual bias, as many questions can be answered without even looking at the visual input. In Appendix A.3, we extend our analysis to NextQA (Xiao et al., 2021).
@@ 3.1 Does Time Matter?
Figure 2: Spatial bias of MVBench video-language benchmark. We show different tasks of the MVBench benchmark and observe that the question can be answered without requiring temporal understanding. For quantitative results see Table 1.
@@
We analyze the spatial bias in MVBench using state-of-the-art image and video-language models such as GPT-4o (OpenAI, 2024), Gemini 1.5 Pro (Gemini, 2024), and Tarsier-34B (Wang et al., 2024a). In Table 1 we focus on four different tasks i) scene transition ii) fine-grained pose iii) fine-grained action and iv) episodic reasoning. To assess whether solving these tasks requires temporal understanding, we compare the performance of such models when receiving only a single random frame, the shuffled videos, and the original videos.
The models receiving only a random frame as input show strong performance across all four tasks in Table 1, surpassing the random baseline. Notably, GPT-4o achieves the highest average performance of 62.8% across the four tasks, nearly matching its video performance with 65.8% and other state-of-the-art video-language models. The lower image performance of Tarsier-34B might stem from its training data composition, which contains five times more video data than image data. These findings are unexpected, as task names like Fine-grained Action suggest a need for temporal understanding. For this fine-grained task, the image-model GPT-4o achieves 49%, which is even slightly better than the state-of-the-art model Tarsier, which scores 48.5%. Similarly, for the other three tasks. Overall, GPT-4o achieves an average accuracy across all 20 tasks of 47.8%, which is 20.5% higher than the random performance of 27.3% on MVBench. This indicates that a large portion of the benchmark is affected by spatial bias.
Additionally, shuffling the videos has minimal impact on the performance of all video-language models with an average difference of 2.3%, indicating that temporal information is not necessary to solve these tasks. Note, as confirmed in Sec. 5.3, the Tarsier model shows a significant drop in performance when videos are shuffled for tasks that require temporal understanding. This problem goes beyond these four tasks as shown in Table 4, Gemini 1.5 Pro and Tarsier achieve an average accuracy across all 20 MVBench tasks of 60.5% and 67.6%, respectively. Shuffling video frames causes a performance drop of only 3.8% and 6.4%, respectively, indicating that the spatial bias affects not only the tasks analyzed in Table 1 but the entire dataset. Additionally, we verify the agreement between the correct responses of Tarsier-34B across modalities: 91.0% between image and video inputs, and 93.9% between video and shuffled video. This confirms that current models heavily rely on spatial biases to solve MVBench.
In Fig. 2 we show examples corresponding to tasks analyzed in Table 1. The example from the left is taken from the Fine-grained Action recognition task implying that temporal understanding is needed. The example on the right is from the Fine-grained Pose task. In both cases, the information provided from any frame is enough to correctly answer the question. In Appendix, in Fig. 14 -20 we show 34 more examples of spatial bias in MVBench.
@@
Table 1: Spatial bias of the MVBench video-language benchmark. Our analysis reveals that temporal understanding is not required for solving these temporal tasks of MVBench as they can be solved with a random frame or shuffled video. Average represents the mean across these four tasks.
Input Fine-grained Action Scene Transition Fine-grained Pose Episodic Reasoning Average
Random 25.0 25.0 25.0 20.0 23.8
Gemini 1.5 Pro image 47.0 78.0 46.5 56.5 57.0
GPT-4o 49.0 84.0 53.0 65.0 62.8
Tarsier-34B 48.5 67.0 22.5 46.0 46.0
Gemini 1.5 Pro video 50.0 93.3 58.5 66.8 67.2
GPT-4o 51.0 83.5 65.5 63.0 65.8
Tarsier-34B 48.5 89.5 64.5 54.5 64.3
Gemini 1.5 Pro shuffle 49.5 90.0 54.5 63.0 64.3
GPT-4o 52.0 84.5 69.0 64.5 67.5
Tarsier-34B 51.0 89.0 56.5 51.5 62.0
@@ 3.2 Does Vision Matter?
Video benchmarks must be designed to prevent questions from being answered solely through common sense reasoning. Modern LLMs possess strong reasoning skills, which can exploit the information within the question and candidate sets in MCQA video language evaluation benchmarks. This creates textual bias, enabling models to answer questions without leveraging the video content.
We analyze the impact of textual bias on MVBench in Table 2. We evaluate the performance of state-of-the-art text-only LLMs, Llama 3 (MetaAI, 2024), and multi-modal LLMs such as Gemini 1.5 Pro (Gemini, 2024), GPT-4o (OpenAI, 2024) and Tarsier (Wang et al., 2024a). Our findings reveal that LLMs can eliminate incompatible candidates easily, greatly outperforming the random baseline. Models using only text achieve competitive results compared to video-language models across these four tasks (Action Count, Unexpected Action, Action Antonym, and Episodic Reasoning). For instance, Gemini 1.5 Pro achieves an average performance of 62.3% using text-only, compared to Tarsier-34B’s 67.4% using videos. Additionally, we verify an 85.3% agreement between Tarsier-34B’s correct text and video responses, confirming its strong reliance on textual biases in MVBench. This goes beyond the four tasks, as Gemini Pro 1.5 achieves an average performance across all 20 tasks of 38.2%, which is 10.9% higher than the random chance baseline of 27.3%. We have identified three key sources of this textual bias.
Bias from LLM-based QA generation. Collecting and manually annotating large datasets for training and evaluation is very costly. Automatic and semi-automatic collection and annotation processes are commonly used (Li et al., 2024b; Mangalam et al., 2024). This includes techniques such as automatic QA pair generation with LLMs. ChatGPT plays a fundamental role in QA generation for 11 of the 20 tasks in the MVBench dataset. However, this introduces unrealistic candidates and QA pairs with excessive information. Fig. 3 presents examples of QA pairs that can be resolved merely with text information. Questions 1 and 2 belong to the Action Antonym task, where an LLM is prompted to generate the antonym of the actual action shown in the video. The answers generated are either unrealistic, as one cannot “remove something into something,” or consistently incorrect, such as “not sure.”
@@
Bias from unbalanced sets. Unbalanced QA sets also hinder a robust evaluation process. For instance, the correct answer for the Action Count task on MVBench is ‘3’ for 90 out of 200 questions, while ‘9’ is only the correct answer for one question. A model with a similar bias might get higher results than random by chance. We have observed in our experiments that some text-only models such as GPT-4o have this bias, predicting ‘3’ for 88 out of 200 samples. This makes GPT-4o perform on par with the best video model with an accuracy of 44.0% and 46.5% respectively.
@@
In Appendix, in Fig. 21 -25 we show 26 more examples of textual bias in MVBench.
Table 2: Textual bias of the MVBench video-language benchmark. Our analysis reveals that vision is not required for solving tasks from MVBench as text-only LLMs score high above the random baseline and nearly on par with video models. Average is with respect to these four tasks.
Input Action Count Unexpected Action Action Antonym Episodic Reasoning Average
Random 33.3 25.0 33.3 20.0 27.9
Llama3 70B text- only 44.5 63.5 74.5 50.5 58.3
Gemini 1.5 Pro 49.0 68.0 85.5 49.0 62.3
GPT-4o 44.0 69.5 57.5 51.5 55.6
Tarsier-34B 37.0 39.5 66.0 44.0 46.6
Gemini 1.5 Pro video 41.2 82.4 64.5 66.8 63.7
GPT-4o 43.5 75.5 72.5 63.0 63.6
Tarsier-34B 46.5 72.0 97.0 54.5 67.4
@@ 4 Open-ended QA to the rescue?
Moreover, as shown in Table 3 open-ended QA does not solve the main issues of MCQA. The performance of text-only models is surprisingly strong; state-of-the-art LLMs can guess the answer solely from the question text for a significant number of questions, even without a candidate list. This includes questions such as Which hand of the person in black wears a watch? or What color is the pants of a person wearing black clothes?, which correct answers are Left hand and Black. The first question can be answered just with prior knowledge as people commonly wear the watch on the left hand, while in the second one, the question gives too much information, the person is wearing black clothes. Similar to the findings for MCQA on spatial bias in Sec. 3.1, when using a single random frame for image-text models such as GPT-4o, performance reaches 60.6% and 46.4%, approaching the video-language model’s 80.3% and 61.6%, respectively. In addition, the performance of Tarsier-34B does not significantly drop—on average less than 3%—when the input videos are shuffled, indicating the low temporal understanding required for solving the benchmarks. These results show that open-ended video-language benchmarks also exhibit strong spatial bias, not requiring temporal understanding to be solved.
@@ 5 TVBench: A temporal video question answering Benchmark
We propose TVBench, a new benchmark specifically created to evaluate temporal understanding in video QA. We adopt a multiple-choice QA approach to prevent the problems of open-ended VQA described in Sec. 4. The main design principles of TVBench are derived from, and address, the problems listed in Sec. 3. Appendix A.2 provides an overview of the tasks, questions, and answers candidates used in our benchmark. We verify our choice of tasks and QA templates in Sec. 5.3 by the performance of multi-modal LLMs with solely a random frame or shuffled videos.
5.1 Developing a temporal video-language benchmark
This section explains the key strategies implemented in TVBench to address the issues identified in Sec. 3 of current video MCQA evaluation benchmarks.
@@ 5.1 Developing a temporal video-language benchmark
MSVD-QA Llama3 70B text 29.1 2.6 51.9 2.8 +22.8
GPT-4o text 29.0 2.5 44.2 2.4 +15.2
GPT-4o image 60.6 3.6 75.3 3.8 +14.7
@@
MSRVTT-QA Lama3 70B text 23.9 2.4 47.8 2.6 +23.9
GPT-4o text 23.3 2.3 42.4 2.3 +19.1
GPT-4o image 34.2 2.7 50.8 2.7 +16.6
@@
ActivityNet-QA Lama3 70B text 25.3 2.6 32.6 1.9 +7.3
GPT-4o text 27.1 2.5 33.9 1.8 +6.8
GPT-4o image 46.4 3.2 56.2 2.9 +9.8
@@
Strategy 1: Define Temporally Hard Answer Candidates. To address Problem 2, it is crucial that the temporal constraints in the question are essential for determining the correct answer. This involves designing time-sensitive questions and selecting temporal challenging answer candidates.
  1. 1.
    We select 10 temporally challenging tasks that require: repetition counting (Action Count), properties of moving objects (Object Shuffle, Object Count, Moving Direction), temporal localization (Action Localization, Unexpected Action), sequential ordering (Action Sequence, Scene Transition, Egocentric Sequence), and distinguishing between similar actions (Action Antonyms).
  2. 2.
    We define hard-answer candidates based on the original annotations to ensure realism and relevance, rather than relying on LLM-generated candidates that are often random and easily disregarded, as seen in MVBench. For example, in the Scene Transition task (Fig. 5), we design a QA template that provides candidates based on the two scenes occurring in the videos for this task, rather than implausible options like “From work to the gym.” Similarly, for the Action Sequence task, we include only two answer candidates corresponding to the actions that actually occurred in the video. More details for the remaining tasks in Appendix A.2.
Strategy 2: Define QA pairs that are not overly informative. Contrary to LLM-based generation, we apply basic templates to mitigate the effect of text-biased QA pairs, mitigating Problem 2.
  1. 1.
    We design QA pairs that are concise and not unnecessarily informative by applying task-specific templates. These templates ensure that the QA pairs lack sufficient information to determine the correct answer purely from text. An example of Unexpected Action is illustrated in Fig. 6. QA pairs require the same level of understanding for the model to identify what is amusing in the video, but without providing additional textual information. Unlike MVBench, the model cannot simply select the only plausible option containing a dog. We use the same candidate sets across tasks like Action Count, Object Count, Object Shuffle, Action Localization, Unexpected Action, and Moving Direction, to ensure balanced datasets with an equal distribution of correct answers, keeping visual complexity while reducing textual bias. Appendix Table 5 provides an overview of all tasks, demonstrating that the QA templates are carefully crafted without unnecessary textual information.
  2. 2.
@@
    Solving the overreliance on world knowledge requires providing questions and candidates that contain only the necessary information, specifically removing factual information that the LLM can exploit. We remove tasks such as Episodic Reasoning, that are based on QA pairs about TV shows or movies.
5.2 Dataset Source
Videos in TVBench are sourced from Perception Test (Patraucean et al., 2024), CLEVRER (Yi et al., 2019), STAR (Wu et al., 2024), MoVQA (Zhang et al., 2023), Charades-STA (Gao et al., 2017), NTU RGB+D (Liu et al., 2019), FunQA (Xie et al., 2023) and CSV (Qian et al., 2022). Overall, TVBench comprises first and third-person perspectives, indoor and outdoor scenes, and real and synthetic data with 2,654 QA pairs among 10 different tasks. QA pairs are generated based on original annotations following the model provided in Appendix A.2 for each task.